# A convergent adaptive finite element stochastic Galerkin method based on multilevel expansions of random fields

### v1

@@ 1.2. Parameter expansions and approximability
with $\beta>1$ and sufficiently small $c>0$, corresponding to the choice $\theta_{0}=1$, $\mathcal{M}=\mathbb{N}$, and $\theta_{j}(x)=cj^{-\beta}\sin(j\pi x)$ for each $j\in\mathcal{M}$. In the limiting case $\beta=1$, the functions $\theta_{j}$ result from the Karhunen-Loève decomposition of the Brownian bridge on $(0,1)$, and with the random coefficients $y\in Y$ distributed according to $\sigma$ in (1.8) one obtains analogous smoothness properties of random draws of $a(y)$.
@@
yields a random field $a$ with very similar features as (1.8), but expanded in terms of the locally supported functions $\theta_{\ell,}$ with $\lvert\operatorname{supp}\theta_{\ell,k}\rvert=2^{-\ell}$. The choice $\alpha=\frac{1}{2}$ corresponds to the well-known Lévy-Ciesielski representation of the Brownian bridge [14].
@@
    for some $\alpha>0$, one has $\lVert\theta_{\mu}\rVert_{L_{\infty}(\Omega)}\lesssim 2^{-\alpha|\mu|}$.
@@ 2. Finite element frames
For newest vertex bisection, we assume the initial triangulation $\hat{\mathcal{T}}_{0}$ to have an edge labelling that is admissible in the sense of [9, Lem. 2.1]; that is, each edge in $\hat{\mathcal{T}}_{0}$ is labelled either 00 or $1$ such that each triangle in $\hat{\mathcal{T}}_{0}$ has exactly two edges labelled $1$ and one edge labelled 00.
@@
There exists $m\in\mathbb{N}_{0}$, $k\in\mathbb{N}_{0}$ and $K\in\mathbb{N}$ such that for all $\mu\in{\mathcal{M}}$, $\theta_{\mu}\in\mathbb{P}_{m}(\hat{\mathcal{T}}_{\lvert\mu\rvert+k})$ and #$\#\{T\in\mathcal{T}_{\lvert\mu\rvert+k}\colon\operatorname{supp}\theta_{\mu}\cap T\neq\emptyset\}\leq K$.
The case of given non-polynomial functions $\theta_{\nu}$ can be treated by replacing them by polynomial approximations. For example, for functions in wavelet-type expansions of Matérn random fields constructed in [4], it is not difficult to obtain suitable efficient polynomial approximations with uniform accuracy that satisfy Assumption 2, see [4, Eq. (139)].
@@
Since for all $K\in\mathcal{T}$ and $E\in\mathcal{E}$, the components $\xi_{K}$ and $\xi_{E}$ are polynomial with degrees bounded by $d_{1}$ and $d_{2}$, respectively, [27, Thm. 3.59] yields
@@
The indices $(\nu,\lambda)$ corresponding to the largest values $\lvert\mathbf{r}_{\nu,\lambda}\rvert$ are selected for refinement, based on a condition analogous to the Dörfler criterion, by a tree thresholding procedure. Subsequently, for each $\nu$, the associated current mesh $\mathcal{T}_{\nu}$ is refined such that all selected frame elements are contained in the resulting finite element space. A similar strategy has been outlined for the refinement of a single finite element mesh in [22]. It is especially in our setting, with interactions between many different meshes in elliptic systems of PDEs, that this technique provides crucial flexbility compared to standard approaches of a posteriori error estimation.
@@ 3. Adaptive solver
In this section, we describe the adaptive solver and analyze its constituent procedures. As noted above, the adaptive scheme is based on Galerkin discretizations on successively refined meshes. Here the way in which the underlying residual error indicators are obtained and used in the mesh refinement differs from previous approaches to adaptive stochastic Galerkin finite element methods. We combine adaptive operator compression in the stochastic degrees of freedom with frame-based spatial refinement indicators. These are subsequently used in a tree-based selection of refinements that in each iteration of the adaptive scheme permits the application of multiple refinement steps within single mesh elements; this latter property is crucial in ensuring error reduction by a uniform factor.
@@ 3.1. Residual estimation
We make use of two particular consequences of Assumption 1: First, for the expansion functions $(\theta_{\mu})_{\mu\in{\mathcal{M}}}$ we have that there exists $C_{1}>0$ such that for all $\ell\geq 0$,
@@
Second, there exists $C_{2}>0$ such that with the $\alpha>0$ in Assumption 1(iii), for all $\ell\geq 0$,
@@
(3.6) #$\#F\leq\sum_{\nu\in F}\#M(\nu)\lesssim\sum_{j=0}^{J}2^{d\ell_{j}}\#F_{j}\lesssim\eta^{-\frac{1}{s}}\bigl{\lVert}\bigl{(}\lVert[v]_{\nu}\rVert_{V}\bigr{)}_{\nu\in{\mathcal{F}}}\bigr{\rVert}_{{\mathcal{A}}^{s}}^{\frac{1}{s}},$
@@ 3.1.1. Approximate dual norm evaluation
$\tilde{\mathcal{T}}_{\nu,\nu^{\prime}}=\{T\in\tilde{\mathcal{T}}_{\nu}\colon\operatorname{supp}[g]_{\nu,\nu^{\prime}}\cap\overline{T}\neq\emptyset\}$
@@
#$\#\bigl{\{}(\tilde{T},\nu)\colon\tilde{T}\in\hat{\mathcal{T}}_{\nu,\nu^{\prime}}\text{ for some $\nu=\nu^{\prime}\pm\mu$, $\lvert\mu\rvert\leq\ell$ and $T\cap\tilde{T}\neq\emptyset$}\bigr{\}}\leq 2M\ell\leq 2M\ell_{\nu^{\prime}}.$
@@
$\sum_{\lvert\mu\rvert\leq\ell_{\nu}}\left(\#\mathcal{T}_{\nu^{\prime}+e_{\mu},\nu^{\prime}}+\#\mathcal{T}_{\nu^{\prime}-e_{\mu},\nu^{\prime}}\right)\leq 2M\ell_{\nu^{\prime}}\#\mathcal{T}_{\nu^{\prime}}+\sum_{\lvert\mu\rvert\leq\ell_{\nu^{\prime}}}2K\lesssim\#\mathcal{T}_{\nu^{\prime}}\ell_{\nu^{\prime}}+2^{d\ell_{\nu^{\prime}}}$
@@
For simplicity, we have assumed the parameter-independent source term $f$ to be piecewise polynomial on the initial triangulation $\hat{\mathcal{T}}_{0}$. However, the above strategy for computing residual indicators can easily be modified to accommodate more general $f$. In particular for $f\in L_{2}(D)$, it suffices to adapt the computation of the single residual component $[r]_{\nu}$ for $\nu=0$ in step (iii) of Algorithm 4, where $f$ can be replaced by an approximation on a sufficiently fine grid according to the current error tolerance. For $f\in H^{-1}(D)$ that are proper functionals, one may instead directly modify step (iv) using the frame coefficients $\langle f,\psi_{\lambda}\rangle$, $\lambda\in\Theta$, and knowledge on their decay properties. Since such modifications are problem-dependent, we do not go into further detail here.

### v2

@@ 1. Introduction
1.1. Motivation: highly sparse representations and convergence of solvers
For a class of ellliptic diffusion problems on a domain $D$, we aim to approximate the mapping from diffusion coefficients $a$ to the corresponding solutions $u$. Here, the first step is the choice of an expansion of $a$ as a function series $a(y)=\theta_{i}+\sum_{i\in\mathbb{N}}y_{i}\theta_{i}$ with $\theta_{i}\in L_{\infty}(D)$, $i\in\mathbb{N}_{0}$, parameterized by a countable family of scalar coefficients $y=(y_{i})_{i\in\mathbb{N}}$. Since the function-valued mapping $y\mapsto u(y)$ is smooth with respect to these parameters, it is natural to approximate it in terms of suitable product polynomials $P_{\nu}(y)=\prod_{i}P_{\nu_{i}}(y_{i})$ with $\deg P_{\nu_{i}}=\nu_{i}$ in the form $u(y)=\sum_{\nu\in F}u_{\nu}P_{\nu}(y)$, where $F$ is a subset of the finitely supported multi-indices in $\mathbb{N}_{0}^{\mathbb{N}}$. Under appropriate assumptions on the decay of $\lVert\theta_{i}\rVert_{L_{\infty}}$ as $i\to\infty$, sparse selections of polynomial degrees $F$ lead to convergence bounds in terms of #$\#F$ without a dependence on the number of active parameters [17, 3].
Note that the coefficients $u_{\nu}$ in such polynomial expansions are still functions on $D$. Their approximability depends on the parameterization chosen for $a$: in particular, when $(\theta_{i})_{i\in\mathbb{N}}$ is a dictionary with wavelet-type multilevel structure, the $u_{\nu}$ each have distinctive localized features that make them particularly amenable to finite element approximations with meshes adapted individually for each $\nu\in F$ [2]. Such adaptive discretizations can be realized based on stochastic Galerkin formulations, that is, joint variational formulations in terms of spatial and parametric variables. With spatial discretizations by adaptive selections from wavelet bases for each $u_{\nu}$, convergent algorithms of optimal computational complexity have been obtained in [6].
However, the common case of adaptive finite element discretizations in space considered in this work poses additional difficulties when allowing independent meshes for each $u_{\nu}$. Since a stochastic Galerkin discretization leads to an elliptic system of PDEs (for #$\#F$ solution components), the associated residuals generally contain edge functionals on various levels of refinement due to the interactions of different meshes. This means that a single step of a standard solve-mark-refine cycle may not necessarily ensure a uniform error reduction [16]. Moreover, since Galerkin orthogonality cannot be ensured practically in this setting, standard residual error estimation techniques are not applicable.
We circumvent both of these issues by an error estimation strategy based on finite element frame coefficients of residuals (used in different contexts in [22, 23]), which does not require Galerkin orthogonality, can naturally handle edge functional terms on different mesh levels, and allows for several refinements of a given element in one step of the adaptive scheme where required. We prove error reduction by a uniform factor in each step of the resulting scheme, which is also known as saturation property. This is the first result of this kind since in the existing works (for example, [18] and [7]) on adaptive stochastic Galerkin methods with independent adaptive meshes for each $u_{\nu}$, this property was instead only assumed to hold. We consequently prove convergence of the derived adaptive method.
We would like to stress that this convergence result by itself does not depend on $(\theta_{i})_{i\in\mathbb{N}}$ in the expansion of $a$ to exhibit a multilevel structure; it could be obtained in the same way for common choices of $\theta_{i}$ that are globally oscillating on $D$. However, we focus on the case of wavelet-like multilevel expansions of $a$ because such expansions are a natural choice, leading to favorable complexity bounds for each component of the adaptive scheme that otherwise are not attainable. Exploiting this type of multilevel structure, we show that the costs for each routine in the adaptive scheme have (up to logarithmic factors) optimal scaling. In our numerical tests, we observe the method to yield approximations converging at the expected optimal rate, with its runtime costs following the same optimal rate.
@@ 1.3. Parameter expansions and approximability
with $\beta>1$ and sufficiently small $c>0$, corresponding to the choice $\theta_{0}=1$, $\mathcal{M}=\mathbb{N}$, and $\theta_{j}(x)=cj^{-\beta}\sin(j\pi x)$ for each $j\in\mathcal{M}$. In the limiting case $\beta=1$, the functions $\theta_{j}$ result from the Karhunen-Loève (KL) decomposition of the Brownian bridge on $(0,1)$, and with the random coefficients $y\in Y$ distributed according to $\sigma$ in (1.8) one obtains analogous smoothness properties of random draws of $a(y)$.
@@
yields a random field $a$ with very similar features as (1.8), but expanded in terms of the locally supported functions $\theta_{\ell,}$ with $\lvert\operatorname{supp}\theta_{\ell,k}\rvert=2^{-\ell}$. The choice $\alpha=\frac{1}{2}$ corresponds to the well-known Lévy-Ciesielski representation of the Brownian bridge [14]. Expansions with similar multilevel structure can also be obtained for more general Gaussian random fields on higher-dimensional domains with Matérn-type or related covariances [4].
@@
    for some $\alpha>0$, one has $\lVert\theta_{\mu}\rVert_{L_{\infty}(D)}\lesssim 2^{-\alpha|\mu|}$.
@@ 1.5. Adaptive scheme and novel contributions
The technique that we use here to obtain a proof of the saturation property does not depend on the wavelet-type multilevel structure of $\{\theta_{\mu}\}_{\mu\in{\mathcal{M}}}$ as given by Assumption 1. It can be used in the same form whenever each $\theta_{\mu}$ can be approximated by piecewise polynomials of fixed degree on a sufficiently fine uniform mesh. It hence can similarly be applied to Karhunen-Loéve-type expansions as in (1.8). Thus, our approach would also lead to a provably convergent method for such more general expansions. The main reason why we focus on expansions of the diffusion coefficient with multilevel structure in what follows is that by exploiting the localization of the $\theta_{\mu}$, we can construct the adaptive solver such that each subroutine has near-optimal computational costs.
@@ 2. Finite element frames
For newest vertex bisection, our starting point is a general initial triangulation $\hat{\mathcal{T}}_{0}$, which we endow with an edge labelling that is admissible in the sense of [9, Lem. 2.1]; that is, each edge in $\hat{\mathcal{T}}_{0}$ is labelled either 00 or $1$ such that each triangle in $\hat{\mathcal{T}}_{0}$ has exactly two edges labelled $1$ and one edge labelled 00. Such a labelling for general conforming two-dimensional triangulations can always be found [9, Lem. 2.1], but requires solving a combinatorial problem. For larger initial meshes, it suffices instead to bisect every triangle twice and take the resulting triangulation as $\hat{\mathcal{T}}_{0}$, where edges that connect two newly created vertices are labelled by 00 and all others by $1$.
@@
Remark 2.1.
Bisection can be used in an analogous manner in higher dimensions, as shown by Stevenson [27]. The essential condition is a guarantee that uniform refinements are again conforming. A construction that guarantees this condition can again be obtained by subdivision; see the appendix of [27]. Although we focus here on $d=2$, using these features of the higher-dimensional analogue of newest vertex bisection, the developments that follow can be extended to $d\geq 3$.
@@
There exists $m\in\mathbb{N}_{0}$, $k\in\mathbb{N}_{0}$ and $K\in\mathbb{N}$ such that for all $\mu\in{\mathcal{M}}$, $\theta_{\mu}\in\mathbb{P}_{m}(\hat{\mathcal{T}}_{\lvert\mu\rvert+k})$ and #$\#\{T\in\hat{\mathcal{T}}_{\lvert\mu\rvert+k}\colon\operatorname{supp}\theta_{\mu}\cap T\neq\emptyset\}\leq K$.
Remark 2.3.
Concerning the role of these assumptions, the following comments are in order:
  1. (a)
    The case of non-polynomial functions $\theta_{\nu}$ can be treated by replacing them by polynomial approximations. Approximability of the $\theta_{\mu}$ of a similar kind are required in any case for quadrature.
  2. (b)
    In its stated form, Assumption 2 applies to typical expansions with multilevel structure such as wavelet-like systems. For example, for functions in wavelet-type expansions of Matérn random fields constructed in [4], it is not difficult to obtain suitable efficient polynomial approximations with uniform accuracy that satisfy Assumption 2, see [4, Eq. (139)]
  3. (c)
    The multilevel structure of $\theta_{\mu}$ implied by Assumption 2 is not required for our main result of uniform error reduction in each step of our adaptive scheme, and thus in particular not for the convergence of the method. As a requirement to obtain this, it would suffice to assume that for a fixed $m$, for each $\mu\in{\mathcal{M}}$ there exists $j(\mu)$ such that $\theta_{\mu}\in\mathbb{P}_{m}(\hat{\mathcal{T}}_{j(\mu)})$. However, the stronger Assumption 2 is crucial for the favorable computational complexity of the adaptive method. We return to this point in Remark 4.5.
  4. (d)
    The same comments apply to $d\geq 3$, where Assumption 2 in its basic form means that each $\theta_{\mu}$ is represented (or approximated) by a piecewise polynomial on a mesh in a hierarchy of uniform refinements and has wavelet-type level-dependent localization.
@@
Since for all $K\in\mathcal{T}$ and $E\in\mathcal{E}$, the components $\xi_{K}$ and $\xi_{E}$ are polynomial with degrees bounded by $d_{1}$ and $d_{2}$, respectively, [28, Thm. 3.59] yields
@@
The indices $(\nu,\lambda)$ corresponding to the largest values $\lvert\mathbf{r}_{\nu,\lambda}\rvert$ are selected for refinement, based on a condition analogous to the Dörfler criterion, by a tree thresholding procedure. Subsequently, for each $\nu$, the associated current mesh $\mathcal{T}_{\nu}$ is refined such that all selected frame elements are contained in the resulting finite element space. A similar strategy has been outlined for the refinement of a single finite element mesh in [22]. It is especially in our setting, with interactions between many different meshes in elliptic systems of PDEs, that this technique provides crucial flexibility compared to standard approaches of a posteriori error estimation.
@@ 3. Adaptive solver
In this section, we describe the adaptive solver and analyze its constituent procedures. As noted above, the adaptive scheme is based on Galerkin discretizations on successively refined meshes. Here, the way in which the underlying residual error indicators are obtained and used in the mesh refinement differs from previous approaches to adaptive stochastic Galerkin finite element methods. We combine adaptive operator compression in the stochastic degrees of freedom with frame-based spatial refinement indicators. These are subsequently used in a tree-based selection of refinements that in each iteration of the adaptive scheme permits the application of multiple refinement steps within single mesh elements; this latter property is crucial in ensuring error reduction by a uniform factor.
The precise procedures for the individual steps in the adaptive scheme are described in this section. They are used by the adaptive solver given in Algorithm 6, which is structured as follows:
  1. (i)
    Finite element frame coefficients of the current residual are approximated by the procedure ResEstimate (Algorithm 4), which in turn has two main steps:
    1. (a)
      First, the semidiscrete operator compression Apply given in Algorithm 1, which selects spatial operator components in the elliptic system of PDEs (1.15) resulting from the stochastic Galerkin formulation. This step can be done analogously to the construction given in [6].
    2. (b)
      Second, finite element frame coefficients of residuals are assembled in Algorithm 4 with the aid of DualNormIndicators (Algorithm 3) and Sum (Algorithm 2). These schemes use localization of the residual contributions according to Lemma 2.4.
  2. (ii)
    In the next step, a selection of frame coefficients is used to produce suitable spatial meshes. This is described in Section 3.2 with the procedures TreeApprox and Mesh, replacing the standard mark-refine step in adaptive finite elements.
  3. (iii)
    A Galerkin solution for the new discretization is computed using GalerkinSolve (Algorithm 5), combining semidiscrete operator compression with standard optimal finite element preconditioning for each spatial component, as described in Section 3.3.
@@ 3.1. Residual estimation
We make use of two particular consequences of 1: First, for the expansion functions $(\theta_{\mu})_{\mu\in{\mathcal{M}}}$ we have that there exists $C_{1}>0$ such that for all $\ell\geq 0$,
@@
Second, there exists $C_{2}>0$ such that with the $\alpha>0$ in 1(iii), for all $\ell\geq 0$,
@@
(3.6) #$\#F\leq\sum_{\nu\in F}\#M(\nu)\lesssim\sum_{i=0}^{I}2^{d\ell_{i}}\#F_{i}\lesssim\eta^{-\frac{1}{s}}\bigl{\lVert}\bigl{(}\lVert[v]_{\nu}\rVert_{V}\bigr{)}_{\nu\in{\mathcal{F}}}\bigr{\rVert}_{{\mathcal{A}}^{s}}^{\frac{1}{s}},$
@@ 3.1.1. Approximate dual norm evaluation
  1. (i)
    Set $L\leftarrow L+1$ and
    $\tilde{\Theta}_{L}=\Big{\{}\lambda^{\prime}\colon\psi_{\lambda^{\prime}}\in\mathcal{R}(\psi_{\lambda})\text{ for $\lambda\in\tilde{\Theta}_{L-1}$ $\wedge$ }$
    $\Big{(}\,\big{(}\xi$ not polynomial on $K$ or $E$ for some $K\in\mathcal{K}(\psi_{\lambda})$, $E\in\mathcal{E}(\psi_{\lambda})\big{)}$
    $\vee$ $\big{(}\operatorname{supp}\psi_{\lambda^{\prime}}\cap K^{\prime}\neq\emptyset$ and $L\leq\operatorname{level}(K^{\prime})+J$ for some $K^{\prime}\in\mathcal{K}\big{)}$
[…]
  2. (ii)
    If $\tilde{\Theta}_{L}\neq\emptyset$ go to (i), else go to (iii);
  3. (iii)
    For $j=L-1,L-2,\dots,0$: For all $\lambda\in\tilde{\Theta}_{j}$ evaluate $\langle\xi,\psi_{\lambda}\rangle$ via
    $\langle\xi,\psi_{\lambda}\rangle=\sum_{\psi_{\lambda^{\prime}}\in\mathcal{R}(\psi_{\lambda})}r_{\lambda,\lambda^{\prime}}\langle\xi,\psi_{\lambda^{\prime}}\rangle\quad\text{if $\mathcal{R}(\psi_{\lambda})\subset\tilde{\Theta}_{j+1}$,}\qquad\text{or by quadrature otherwise;}$;
  4. (iv)
    Set $\Theta^{+}=\bigcup_{j=0}^{L-1}\tilde{\Theta}_{j}$ and return $\bigl{(}\Theta^{+},\,\bigl{(}\langle\xi,\psi_{\lambda}\rangle\bigr{)}_{\lambda\in\Theta^{+}}\bigr{)}$.
@@
$\tilde{\mathcal{T}}_{\nu,\nu^{\prime}}=\{T\in\tilde{\mathcal{T}}_{\nu}\colon\operatorname{supp}[g]_{\nu,\nu^{\prime}}\cap T\neq\emptyset\}$
@@
#$\#\bigl{\{}(\tilde{T},\nu)\colon\tilde{T}\in\hat{\mathcal{T}}_{\nu,\nu^{\prime}}\text{ for some $\nu=\nu^{\prime}\pm e_{\mu}$, $\lvert\mu\rvert\leq\ell$ and$T\cap\tilde{T}\neq\emptyset$}\bigr{\}}\leq 2M\ell\leq 2M\ell_{\nu^{\prime}}.$
@@
(3.14) $\sum_{\lvert\mu\rvert\leq\ell_{\nu}}\left(\#\mathcal{T}_{\nu^{\prime}+e_{\mu},\nu^{\prime}}+\#\mathcal{T}_{\nu^{\prime}-e_{\mu},\nu^{\prime}}\right)\leq 2M\ell_{\nu^{\prime}}\#\mathcal{T}_{\nu^{\prime}}+\sum_{\lvert\mu\rvert\leq\ell_{\nu^{\prime}}}2K\lesssim\#\mathcal{T}_{\nu^{\prime}}\ell_{\nu^{\prime}}+2^{d\ell_{\nu^{\prime}}}$
@@
For simplicity, as stated in the beginning of Section 2, so far we have assumed the parameter-independent source term $f$ to be piecewise polynomial on the initial triangulation $\hat{\mathcal{T}}_{0}$. However, the above procedure for computing residual indicators can easily be modified to accommodate more general $f$. There are several possible ways of achieving this by straightforward adaptation of standard strategies. To avoid additional technicalities, we restrict ourselves to briefly outlining two options:
  1. (a)
    A general approach is to use an additional routine $\text{\sc Approx}_{f}(\eta)$ that returns an approximation $\tilde{f}$ of $f$ satisfying $\lVert\tilde{f}-f\rVert_{H^{-1}(D)}\leq\eta$ and is of the form as $\xi$ in Lemma 2.4 for some triangulation. This approximation can be used in step (iii) of Algorithm 4. Then we can guarantee the result of 3.7 again with the help of Lemma 2.4. A very similar strategy was used in [26, Sec. 6]. The construction of $\text{\sc Approx}_{f}(\eta)$ depends on the particular features of $f$: if $f$ is in $L_{2}(D)$, such a routine can be obtained by $L_{2}$-projections onto a suitable space of piecewise polynomials; for proper functionals $f\in H^{-1}(D)$, such approximations are more problem-specific, as considered in detail in [16].
  2. (b)
    We can also more directly use knowledge of the coefficients $\langle f,\psi_{\lambda}\rangle$, $\lambda\in\Theta$ and their decay to provide estimates
    $\Bigl{(}\sum_{\lambda\in\Theta\setminus\Theta^{f}_{\eta}}\lvert\langle f,\psi_{\lambda}\rangle\rvert^{2}\Bigr{)}^{\frac{1}{2}}\leq\eta$
    for certain finite subsets $\Theta^{f}_{\eta}\subset\Theta$. These can be used as input in step (iv) of Algorithm 4.